Let PP represent the profit (or loss), in thousands of dollars, for a certain company tt years after 1960. Based on the data shown below, the best-fit line is given to be

P=−2.42t+37.93P=-2.42t+37.93

tt PP
1 35.24
2 32.87
3 33.1
4 27.23
5 25.26
6 23.39
7 22.32
8 17.55
9 15.08
10 11.81
11 12.54
12 9.07
13 7.5



Use the given best-fit line to estimate the profit in 1968.

The esitmated profit (or loss) in 1980 is predicted to be $ (thousands of dollars).
Answer #1

For 1968, x = 1968 - 1960 = 8

So,

Predicted profit = -2.42*8 + 37.93 = 18.57

For 1980, x = 1980 - 1960 = 20

So,

Predicted profit/ loss = -2.42*20 + 37.93 = -10.47